OpenStudy (anonymous):

Find the sum. 4x2 + 7x2 = 10x2 10x4 11x2 11x4

OpenStudy (tyteen4a03):

Hint: Those terms are like terms. More obvious hint: 7a^2 + 3a^2 = 10a^2.

OpenStudy (kendricklamar2014):

@Haksorin can you tell me what 7+4 = ?

OpenStudy (anonymous):

Twas D

OpenStudy (kendricklamar2014):

Yes, D is the answer because 7+4 = 11, then you add the x^2 to make the answer: 11x^2

OpenStudy (tyteen4a03):

@Haksorin @KendrickLamar2014 No. A like term consists of two parts: The coefficient (the number you see before the variable) and the variable itself, including any exponents (such as \(x^2\) or \(x^7\). In this case, you are adding \(4x^2 + 7x^2\). The variable is \(x^2\) and must remain unchanged.

OpenStudy (kendricklamar2014):

You got the answer wrong @tyteen4a03 You replaced 4 with 3. Im just saying that 7x^2+4x^2 = 11x^2

OpenStudy (tyteen4a03):

@KendrickLamar2014 Answer D is \(11x^4\), not \(11x^2\).

OpenStudy (kendricklamar2014):

Typing error Instead of putting 4 I put 2 :P